OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A.y+5=6(x-1) B.y+1=6(x-5) C.y-5=6(x+1) D.y-1=6(x+5)

OpenStudy (anonymous):

x1 = 5, y1 = -1 Equation: (y - y1) = m(x - x1) => y+1 = 6(x-5)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thanks